OpenStudy (anonymous):

Find the mass in grams of 2.00*10^23 molecules of F2.

OpenStudy (anonymous):

\[2.00 \times 10^{23} molecules F _{2} \times \frac{ 1mol F _{2} }{ 6.023 \times 10^{23} molecules F _{2} } \times \frac{ 38.00 g F _{2} }{ 1mol F _{2} }\]

OpenStudy (anonymous):

1molF2, would that be the atomic mass or Fluorine times 2?

OpenStudy (anonymous):

Yep. 1 mole is atomic mass.

OpenStudy (anonymous):

would the answer be 1.26*10^47?

OpenStudy (anonymous):

If you calculated it right, it would be.

OpenStudy (anonymous):

Okay, thank you very much!!

OpenStudy (anonymous):

Doenst look like it to me. It should be less.

OpenStudy (anonymous):

It should be way way less. The grams should be less than 38 grams.

OpenStudy (anonymous):

on my calculator I typed in 2.00*10^23 x 37.996/6.022*10^23 x 38.00/37.996 = 1.26203919*10^47

OpenStudy (anonymous):

Where are you getting those numbers? Use my dimensional analysis. As you can see the given unit is canceled out and the mass unit (grams) is left. No need to add anything.

OpenStudy (anonymous):

I did use it. I replace 1molF2 w/ 37.996, thats the only thing that is different!

OpenStudy (anonymous):

I got that by doing: (2*18.998) = 37.996

OpenStudy (anonymous):

Close enough to be 38.00 gram. Thats why I put it on top of mole. 1 mole F2=38.00 grams F2 You dont replace anything in this conversion. You can but its not necessary. Just makes everything longer. And you are not entering it in the calculator right because of the unnecessary steps. (2.00 x 10^23/6.023 x 10^23) x 38.00 This is just basic operations.

OpenStudy (anonymous):

okay well thank you

OpenStudy (anonymous):

No problem!